1.      A college administrator is interested in determining the proportion of students who receive some sort of financial aid. Instead of examining the records for all students, the administrator randomly draws 120 students and finds that 67 of them are receiving financial aid. Use a 90% confidence interval to estimate the true proportion of students who receive financial aid. (1)

 

2.      In estimating the population mean with the population standard deviation unknown, if the sample size is 17, how many degrees of freedom will there be? (1)

1
Expert's answer
2021-03-29T07:28:58-0400

1) "p=\frac{67}{120}=0.5583."

"90\%CI=(\frac{67}{120}-1.645\sqrt{\frac{0,5583(1-0.5583)}{120}},\frac{67}{120}+1.645\sqrt{\frac{0,5583(1-0.5583)}{120}})="

"=(0.4838,0.6329)."


2) "df=17-1=16."